OpenStudy (anonymous):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. b over fifteen minus one equals twenty-nine Step 1: b over fifteen minus one plus one equals twenty-nine plus one Step 2: b over fifteen equals thirty Step 3: b over fifteen divided by fifteen equals thirty divided by fifteen Step 4: b equals two

OpenStudy (anonymous):

b/15 - 1 = 29 b/15 = 30 (b/15)/15 = 30/15 b/225 = 2 . . .

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Seems like Yuri made a mistake between steps 3 and 4 where he divided by 15 instead of multiplying by 15.

OpenStudy (anonymous):

what yuou mean